Find n, given that we wish to estimate Mean to within 10 units, with 95% confidence, and assuming that Standard Deviation=100.


1
Expert's answer
2021-03-18T02:29:54-0400

"\text{Given that}\\ \sigma=100,E=10,\alpha=0.05,\\
\therefore Z_{\frac{\alpha}{2}}=Z_{0. 025}=1.96\\n=\frac{Z_{\frac{\alpha} {2}}^{2}\times \sigma^{2}} {E^{2}}\\
n=\frac{1.96^{2}\times 100^{2}} {10^{2}} =384.16\approx 385"